OpenStudy (anonymous):

2^x * 4^x / 8^x - 1 Express in simplest form with a prime number base. Please Explain.

OpenStudy (campbell_st):

ok... so just checking its \[\frac{2^x \times 4^x}{8^{x - 1}}\]

OpenStudy (anonymous):

Yeah

OpenStudy (campbell_st):

ok... well you need to know a few things about 4 and 8 \[\frac{2^x \times (2^2)^x}{(2^3)^{x - 1}}\] now all the terms are in the same base... so the index laws you need are 1. Power of a power... multiply the powers \[(x^a)^b = x ^{a \times b}\] 2. multiplying terms with the same base, add the powers \[x^a \times x^b = x^{a + b}\] 3. dividing the same base, subtract the powers \[\frac{x^a}{x^b} = x^{a - b}\] hope it all helps..

OpenStudy (anonymous):

Thank you :)